# Life Sequence Transformer: Generative Modelling for Counterfactual Simulation

### v1

@@ A.1 Event Study
In such cases, researchers often turn to an event study design [Jacobson et al., 1993], a generalisation of DiD that exploits variation in treatment timing. Event studies are especially useful when all units are eventually treated at different times. The method compares changes in outcomes before and after treatment for each unit, leveraging variation across units treated at different points in time. In the childbirth example, this allows for estimating treatment effects by aligning individuals around the “event” (birth of a child) and comparing pre- and post-treatment trajectories. Like DiD, event studies rely on a version of the parallel trends assumption—that, absent the event, the trajectory of the outcome would have continued as observed in the pre-treatment period.
A.2 Synthetic Control
@@
This method was originally proposed by Abadie et al. [2010] to study the economic impact of California’s 1988 tobacco control program. Rather than comparing California to a single state, SCM used a weighted average of other U.S. states to construct a “synthetic California” that closely mirrored its economic indicators before the intervention. Any divergence between the treated unit and its synthetic control after treatment can then be attributed to the intervention, under the assumption that the synthetic control approximates the counterfactual trajectory the treated unit would have followed in the absence of treatment.
@@ Appendix B Work History Italian Panel Data
The Work History Italian Panel (WHIP) [Leombruni et al., 2010, Bena et al., 2012] consists of a systematic sample of 1 in 15 individuals drawn from the INPS database based on their date of birth. This database allows us to track the employment histories of sampled individuals, including all periods of employment recorded by INPS, retirement, and any time spent receiving social security benefits, such as unemployment subsidies or maternity leave. The WHIP currently covers the period from 1985 to 2019.
@@
    Table 2: Sample selection. Each row in the table corresponds to a selection criterion. The second column reports the percentage of individuals discarded, regardless of the other selection criteria
@@ 45.11 Trade in passenger cars and light motor vehicles
The last set of features provides details about the individual’s income and its nature. (7) Labour force status is specific to WHIP, i.e. it does not follow any international classification system. It describes the type of income, i.e. in what context the income is associated with a person. Examples of these statuses include: "Employee", "Pensioner", "Self-employed", "Unemployed", encompassing work-related and other socioeconomic statuses. "Employee", "Self-Employed, or "Para-Subordinate" (e.g. sub-contractors, consultants, people with work contracts but not all the benefits) have one additional modifying information, the (8) Working title which indicates the qualification in the job and may fall into one of 11 categories: in training, Workman, Clerk, Junior Manager, Senior Manager, technical-administrative, Sportsman, Artist, Voucher, Collaborator, or Professional.
For "Employee" records only, we have four additional modifiers: (9) work intensity, (10) maternity leave intensity, (11) sick leave intensity, and whether the work is (12) part-time or full-time. The three intensities describe the average number of weeks in a month when an employee was working, on maternity leave, or sick. We derive the intensity by dividing the overall number of weeks spent in a given category (e.g. working) by the total number of months the employment relationship lasted in a given year. Therefore, each record can take a maximum value of around 4.3 (52 weeks divided by twelve), and a minimum value of 0. Notably, maternity leave is only recorded for female employees. Finally, we discretise the intensity into five categories based on the rounded value. If, for example, the intensity work is zero, we encode it as <WRKINT_S0>, if the intensity is less than 1, we encode it as <WRKINT_S1>, for intensities greater than 4, they are encoded into the category <WRKINT_S4+>.
The final attributes are the (13) month the event started, its (14) duration in months within a year, and (15) income during the event. The nature of income differs depending on the labour force status. For example, for "Employees", "Self-employed", "Para-Subordinate", and "Rentiers", it is the reported income from which the INPS calculates social security contributions. For "Unemployed" and "Pensioner", the income is the amount that INPS provides to the individual. The raw version of the income is a yearly value; we divide it by the duration in months of a given labour force status within a year and derive a monthly income. We further adjust it for inflation, deriving a real monthly income, and discretise this value into 100 quantiles.
@@
Each life sequence is represented as a list of discrete tokens, encoding time-indexed events and their characteristics. Sequences are capped at a maximum of 1,560 tokens. Any sequence shorter is padded with a padding token, and any sequence longer is truncated starting from the earliest years (i.e. from the start of the sequence), ensuring that all input tensors are of consistent shape for training and inference. Longer sequences are truncated by whole years so as not to lose the temporal structure of the sequence (i.e. all tokens in a year are deleted until the sequence is at most 1,560 tokens long).
@@
i.e. a female born in January of 1942 in the Center region of Italy. Assume that in the first year of measurement (1990), she worked full time during the whole year, then had a year of no recorded activity, and then got another job in a different province for the entire following year:
@@ Appendix C Model architecture
Transformers are a class of neural network architectures initially developed for natural language processing tasks, such as machine translation or text summarization [Vaswani et al., 2017], but are now being applied to a wide variety of sequence types from different domains, including protein discovery Consens et al. [2025], health predictions [Yang et al., 2023], music generation Agostinelli et al. [2023]. Their strength lies in the self-attention mechanism, which enables them to model long-range dependencies in sequential input.
@@
One of the most influential encoder-only Transformer models is BERT (Bidirectional Encoder Representations from Transformers), introduced by Devlin et al. [2019]. BERT is trained using a masked language modelling (MLM) objective, where a subset of tokens in the input sequence is randomly masked, and the model learns to predict them using bidirectional context—i.e., both left and right tokens. This bidirectional attention mechanism enables BERT to build rich, context-aware representations of entire sequences, making it highly effective for discriminative tasks such as text classification, question answering, and semantic similarity. Since BERT processes the entire input at once, it excels at tasks that require a holistic understanding of the input sequence, rather than step-by-step prediction. An encoder-only version of our model is similar to Savcisens et al. [2024]’s Life2Vec, which is inspired by the BERT architecture and training objective.
@@
In contrast, GPT (Generative Pre-trained Transformer) exemplifies a decoder-only architecture. Introduced by Radford et al. [2018], GPT is trained with an auto-regressive objective, predicting the next token in a sequence based solely on the preceding tokens. GPT uses causal (masked) self-attention, which restricts each token’s attention to preceding positions in the sequence, preventing information from "leaking" from future tokens. Unlike BERT, GPT does not explicitly mask input tokens, but the causal attention mechanism ensures that generation proceeds in a left-to-right manner. This makes the GPT architecture particularly well-suited for generative tasks, such as language generation, sequential modelling, and, in our case, simulating counterfactual life sequences. Figure 6 illustrates the decoder-only architecture we train in this paper.
@@
Figure 6: Overall architecture of the decoder-only model. Each tokenised life sequence is passed through an embedding block and a stack of decoder blocks. Each token—consisting of concept, age, and year information—is embedded using three components: concept embeddings retrieved from a learned lookup matrix $\mathcal{E}_{\mathcal{V}}$ (normalised by subtracting its mean), and temporal features (age and year) encoded via Time2Vec. These are combined with learnable positional embeddings to form the input token representations. The resulting sequence is processed through multiple decoder blocks, each structured to preserve the autoregressive nature of the task. Each block starts with RMS normalisation, followed by a masked multi-head attention layer (using the Performer mechanism with local and global heads), ensuring that each token only attends to preceding tokens. The output of the attention layer is combined with the input via a residual connection. This is followed by another RMS normalisation and a position-wise feed-forward layer, whose output is again added through a residual connection. Finally, the contextualised output representations are passed through a linear projection head and to a softmax function that assigns a probability to each token in the vocabulary. The predicted token is the one with the highest probability score.
Between the development of encoder-only models like BERT and decoder-only models like GPT, another approach combined the strengths of both: the encoder–decoder architecture. Popularised by models such as T5 (Text-to-Text Transfer Transformer), this architecture was initially inspired by earlier sequence-to-sequence models used in machine translation [Raffel et al., 2020]. In this setup, an encoder processes an input sequence—such as a question, a prompt, or a contextual situation—and produces a latent representation that is then consumed by a decoder to generate a corresponding output sequence, such as an answer, a summary, or a translated sentence. The decoder uses cross-attention to condition its generation on the full encoded representation, effectively enabling the model to incorporate rich contextual information while generating responses in an autoregressive manner.
Encoder–decoder models are especially well-suited to tasks where there is a clear distinction between input and output roles, and where conditioning on the full input is crucial—examples include question answering, summarisation, and program synthesis. While this approach is more computationally demanding—requiring separate encoder and decoder stacks and more memory for cross-attention—it allows for greater flexibility and generality, especially in multi-modal or structured input scenarios. In our context of life sequence modelling, encoder–decoder architectures are particularly appealing when we wish to condition the prediction of future life trajectories on a fixed set of background covariates (e.g., age, sex, place of birth), which can be encoded once and used to guide generation through the decoder.
Figure 7 illustrates the encoder-decoder architecture we develop and train in this paper. The encoder (right side of the figure) processes a short, unordered sequence of concept tokens representing background characteristics (e.g., gender, education). Since these inputs are unordered, positional embeddings are omitted. Each token is passed through an embedding layer, then into a standard multi-head self-attention layer. RMS Normalisation is applied before both the attention and feedforward layers. The output of the attention layer is added back to the input via a residual connection. The final output is a fixed-length contextualised representation that conditions the decoder via cross-attention. The decoder (left side) follows an autoregressive design. Each concept token in a sequence (including age and year information) is passed through an embedding block that merges concept embeddings, age and year embeddings (via time2vec), and rotary positional encodings (RoPE) to capture temporal structure. The concept embeddings are drawn from a lookup matrix $\mathcal{E}_{\mathcal{V}}$, with the sequence mean subtracted to stabilise learning.
The decoder comprises stacked blocks, each with three components: RMS Normalisation, a causal (masked) self-attention layer using Performer Attention, and a feedforward layer. In the first two decoder blocks, an additional cross-attention mechanism allows the decoder to attend to the encoder outputs. RMS Normalisation precedes attention and feedforward components, and residual connections are applied throughout. During training, the model is optimised for a next-token prediction objective, learning to generate plausible and coherent life sequences conditioned on prior events and background features.
Figure 7: Overall architecture of the encoder-decoder model.
@@ Appendix D Experimental details
Retirement Timing
As a first causal benchmarking exercise, we evaluate the model’s ability to replicate the well-documented finding that month of birth influences retirement timing [Ardito et al., 2020]. Specifically, individuals born earlier in the calendar year tend to retire slightly later, due to institutional and lifecycle timing mechanisms.
To assess this, we select from the test set all male individuals born between 1940 and 1950, restricting the sample to those born in either January or December. To ensure a clean employment trajectory, we include only those who have been continuously and exclusively employed in standard employment, excluding individuals who were ever self-employed, para-subordinate, unemployed, or employed under special regimes (e.g., artists, athletes, or voucher-based contracts). We further restrict the sample to those whose first pension spell occurs between ages 55 and 68, and who exhibit no long gaps in their administrative records.
To reduce noise from incomplete information, we also exclude individuals with missing values in key attributes such as work province, economic sector (ATECO), income, and job title. After applying these criteria, the final evaluation set comprises 1,010 individuals.
For each of these individuals, we provide the model with their observed sequence up to the year before retirement, and task it with generating their life sequences forward until 2015. We therefore analyse whether the model both captures the correct retirement timing and the expected shift between those born in January and December.
Child Penalty
The second exercise evaluates the model’s capacity to reproduce the income loss experienced by women following childbirth, a well-established finding in the causal literature on gender inequality in the labour market [Kleven et al., 2019, Casarico and Lattanzio, 2023].
From the test set, we selected a sample of women who had only held standard employment contracts throughout their recorded work histories. We identify mothers as those who experienced at least one maternity leave spell, defined by a maternity level intensity greater than one. To ensure sample homogeneity, we exclude individuals who were ever classified as managers or professional athletes, as these careers may follow atypical trajectories.
Further, we restrict the sample to women employed in southern Italian provinces at the time of childbirth, as here labour market dynamics tend to be more constrained [Carta et al., 2024]. To ensure high data quality, we retain only those individuals with less than 1% missing data across key variables: economic sector (ATECO), work province, income, and job title.
To allow for sufficient observation both before and after childbirth, we focus on women whose first childbirth occurs between the 5th and 20th year of their observed life course. This ensures a meaningful pre-treatment period for conditioning and a post-treatment period for simulation.
The final sample comprises 631 mothers. For each, we provide the model with their complete life sequence up to the year of childbirth, and generate the trajectory from that point until the end of the observation window. This allows us to evaluate whether the model captures the expected post-maternity income trend.
Mobility allowance duration

### v2

@@ A.1 Event Study
In such cases, researchers often turn to an event study design (Jacobson et al., 1993), a generalisation of DiD that exploits variation in treatment timing. Event studies are especially useful when all units are eventually treated at different times. The method compares changes in outcomes before and after treatment for each unit, leveraging variation across units treated at different points in time. In the childbirth example, this allows for estimating treatment effects by aligning individuals around the “event” (birth of a child) and comparing pre- and post-treatment trajectories. Like DiD, event studies rely on a version of the parallel trends assumption—that, absent the event/treatment, the trend of the outcome would have continued as observed in the pre-treatment period.
A.2 Propensity Score
While Difference-in-Differences (DiD) relies on the parallel trends assumption, in many empirical settings, treatment and control groups differ substantially in observable characteristics, making the assumption less credible. One strategy to mitigate this concern is to combine DiD with Propensity Score Matching (PSM), a method introduced by Rosenbaum and Rubin (1983). The core idea of PSM is to construct a control group that is statistically similar to the treated group in terms of observed covariates, by matching treated and untreated units with comparable estimated probabilities of treatment (the “propensity score”).
When PSM is combined with DiD, the matching procedure reduces bias from differences in observable characteristics, while the DiD framework accounts for unobserved time-invariant heterogeneity. This hybrid approach is particularly valuable when treatment assignment is not random and when untreated units differ systematically from treated ones. By matching first and then applying DiD, researchers ensure that the parallel trends assumption is more plausible for the reweighted or matched sample.
A.3 Synthetic Control
@@
This method was originally proposed by Abadie et al. (2010) to study the economic impact of California’s 1988 tobacco control program. Rather than comparing California to a single state, SCM used a weighted average of other U.S. states to construct a “synthetic California” that closely mirrored its economic indicators before the intervention. Any divergence between the treated unit and its synthetic control after treatment can then be attributed to the intervention, under the assumption that the synthetic control approximates the counterfactual trajectory the treated unit would have followed in the absence of treatment.
@@ Appendix B Work History Italian Panel Data
The Work History Italian Panel (WHIP) (Leombruni et al., 2010; Bena et al., 2012) consists of a systematic sample of 1 in 15 individuals drawn from the INPS database based on their date of birth. This database allows us to track the employment histories of sampled individuals, including all periods of employment recorded by INPS, retirement, and any time spent receiving social security benefits, such as unemployment subsidies or maternity leave. The WHIP currently covers the period from 1985 to 2019.
@@
    Table 3: Sample selection. Each row in the table corresponds to a selection criterion. The second column reports the percentage of individuals discarded, regardless of the other selection criteria.
@@ 45.11 Trade in passenger cars and light motor vehicles
The last set of features provides details about the individual’s income and its nature. (7) Labour force status is specific to WHIP, i.e., it does not follow any international classification system. It describes the type of income, i.e., in what context the income is associated with a person. Examples of these statuses include: Employee, Pensioner, Self-employed, Unemployed, encompassing work-related and other socioeconomic statuses. Employee, Self-Employed, or Para-Subordinate (e.g., sub-contractors, consultants, people with work contracts but not all the benefits) have one additional modifying information, the (8) Working title which indicates the qualification in the job and may fall into one of 11 categories: in training, Workman, Clerk, Junior Manager, Senior Manager, technical-administrative, Sportsman, Artist, Voucher, Collaborator, or Professional.
For Employee records only, we have four additional modifiers: (9) work intensity, (10) maternity leave intensity, (11) sick leave intensity, and whether the work is (12) part-time or full-time. The three intensities describe the average number of weeks in a month when an employee was working, on maternity leave, or sick. We derive the intensity by dividing the overall number of weeks spent in a given category (e.g., working) by the total number of months the employment relationship lasted in a given year. Therefore, each record can take a maximum value of around 4.3 (52 weeks divided by twelve), and a minimum value of 0. Notably, maternity leave is only recorded for female employees. Finally, we discretise the intensity into five categories based on the rounded value. If, for example, the intensity work is zero, we encode it as <WRKINT_S0>, if the intensity is less than 1, we encode it as <WRKINT_S1>, for intensities greater than 4, they are encoded into the category <WRKINT_S4+>.
The final attributes are the (13) month the event started, its (14) duration in months within a year, and (15) income during the event. The nature of income differs depending on the labour force status. For example, for Employees, Self-employed, Para-Subordinate, and Rentiers, it is the reported income from which the INPS calculates social security contributions. For Unemployed and Pensioner, the income is the amount that INPS provides to the individual. The raw version of the income is a yearly value; we divide it by the duration in months of a given labour force status within a year and derive a monthly income. We further adjust it for inflation, deriving a real monthly income, and discretise this value into 100 quantiles.
@@
Each life sequence is represented as a list of discrete tokens, encoding time-indexed events and their characteristics. Sequences are capped at a maximum of 1,560 tokens. Any sequence shorter is padded with a padding token, and any sequence longer is truncated starting from the earliest years (i.e., from the start of the sequence), ensuring that all input tensors are of consistent shape for training and inference. Longer sequences are truncated by whole years so as not to lose the temporal structure of the sequence (i.e., all tokens in a year are deleted until the sequence is at most 1,560 tokens long).
@@
i.e., a female born in January of 1942 in the Center region of Italy. Assume that in the first year of measurement (1990), she worked full time during the whole year, then had a year of no recorded activity, and then got another job in a different province for the entire following year:
@@ Appendix C Model architecture
Transformers are a class of neural network architectures initially developed for natural language processing tasks, such as machine translation or text summarization (Vaswani et al., 2017), but are now being applied to a wide variety of sequence types from different domains, including protein discovery Consens et al. (2025), health predictions (Yang et al., 2023), music generation Agostinelli et al. (2023). Their strength lies in the self-attention mechanism, which enables them to model long-range dependencies in sequential input.
@@
One of the most influential encoder-only Transformer models is BERT (Bidirectional Encoder Representations from Transformers), introduced by Devlin et al. (2019). BERT is trained using a masked language modelling (MLM) objective, where a subset of tokens in the input sequence is randomly masked, and the model learns to predict them using bidirectional context, i.e., both left and right tokens. This bidirectional attention mechanism enables BERT to build rich, context-aware representations of entire sequences, making it highly effective for discriminative tasks such as text classification, question answering, and semantic similarity. Since BERT processes the entire input at once, it excels at tasks that require a holistic understanding of the input sequence, rather than step-by-step prediction. An encoder-only version of our model is similar to Savcisens et al. (2024)’s Life2Vec, which is inspired by the BERT architecture and training objective.
@@
In contrast, GPT (Generative Pre-trained Transformer) exemplifies a decoder-only architecture. Introduced by Radford et al. (2018), GPT is trained with an auto-regressive objective, predicting the next token in a sequence based solely on the preceding tokens. GPT uses causal (masked) self-attention, which restricts each token’s attention to preceding positions in the sequence, preventing information from “leaking” from future tokens. Unlike BERT, GPT does not explicitly mask input tokens, but the causal attention mechanism ensures that generation proceeds in a left-to-right manner. This makes the GPT architecture particularly well-suited for generative tasks, such as language generation, sequential modelling, and, in our case, simulating counterfactual life sequences. Figure 7 illustrates the decoder-only architecture we train in this paper.
@@
Figure 7: Overall architecture of our decoder-only model. Each tokenised life sequence is passed through an embedding block and a stack of decoder blocks. Each token, consisting of concept, age, and year information, is embedded using three components: concept embeddings retrieved from a learned lookup matrix $\mathcal{E}_{\mathcal{V}}$ (normalised by subtracting its mean), and temporal features (age and year) encoded via Time2Vec. These are combined with learnable positional embeddings to form the input token representations. The resulting sequence is processed through multiple decoder blocks, each structured to preserve the autoregressive nature of the task. Each block starts with RMS normalisation, followed by a masked multi-head attention layer (using the Performer mechanism with local and global heads), ensuring that each token only attends to preceding tokens. The output of the attention layer is combined with the input via a residual connection. This is followed by another RMS normalisation and a position-wise feed-forward layer, whose output is again added through a residual connection. Finally, the contextualised output representations are passed through a linear projection head and to a softmax function that assigns a probability to each token in the vocabulary. The predicted token is the one with the highest probability score.
@@ Appendix D Experimental details
This section outlines the experimental setup, notation, and estimands, as well as the process of converting generated simulations into treatment effects that are directly comparable to their empirical counterparts. Our goal is to place generative model-based evaluation on the same footing as standard causal inference designs in applied economics. We build on the event study and propensity-score frameworks described in Appendix A.1A.2, adapting them to the context of sequences generated from our model.
Common setup and notation
Let $i$ index selected individuals in the test set. Each empirical exercise is framed in terms of a binary treatment $T_{i}$ and a scalar outcome $Y_{i}$. The central estimand is the Average Treatment Effect (ATE), defined as the expected difference in outcomes between treated and untreated/control units:
$\text{ATE}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)],$
where $Y_{i}(1)$ and $Y_{i}(0)$ denote the potential outcomes for individual $i$ under treatment and control, respectively. In practice, we never observe both potential outcomes for the same person; however, under appropriate research designs (such as random assignment, natural experiments, or valid identification strategies), we can estimate the ATE as the difference in mean outcomes between the treated and control groups. In our experiment, these quantities are:
  • Pension: $T_{i}=1$ if individual $i$ was born in January, $T_{i}=0$ if born in December. The outcome $Y_{i}$ is the age at first pension claim, measured in months.
  • Unemployment: $T_{i}=1$ if individual $i$ experienced their first unemployment spell at or after 40 years of age (480 months), $T_{i}=0$ otherwise. The outcome $Y_{i}$ is the number of months spent in the first unemployment spell. The estimand is local, defined within a symmetric window $h$ around the 480-month cutoff.
  • Maternity: $T_{i}=1$ if individual $i$ experiences at least one maternity leave episode in her sequence, $T_{i}=0$ otherwise. The outcome $Y_{i}$ is the average post-maternity annual earnings.
For maternity, treatment is endogenous rather than rule-based. As such, the definition of a control group is non-trivial. Thus, we employ an event study design and a propensity score matching approach (Section A.1) to align careers around childbirth events and identify a suitable control group of non-mothers that matched he pre-birth characteristics of the mothers. To make results comparable across experiments, we summarise the estimated time-path into a single number: the mean effect over the birth year and the first ten post-birth years (see Section D.2).
From simulations to outcomes
For every person $i$ we generate $K$ continuations from a fixed cutoff of their observed history. This cutoff point corresponds to the moment just before the outcome of interest takes place:
  • Pension: one year before observed retirement (denoted 1y in tables) or four years before retirement (4y). An example of the former setup is shown in Figure 8.
  • Unemployment: the year of the first observed unemployment spell.
  • Maternity: the year of the first observed maternity leave (offset–0), as well as one year prior (offset–1), to allow counterfactual generation without maternity.
(a) Original sequence
(b) Sequence Cut
(c) Final input sequence
Figure 8: Example of the simulation setup: (a) shows the original work sequence, (b) the sequence cut applied one year before retirement occurs, and (c) the resulting input sequence used for simulation.
Each simulation $k$ yields a synthetic outcome $\tilde{Y}_{ik}$. We compute the person-level Monte Carlo mean
$Y_{i}^{\text{sim}}\;=\;\frac{1}{K}\sum_{k=1}^{K}\tilde{Y}_{ik},$
and then use $\{Y_{i}^{\text{sim}}\}_{i=1}^{n}$ to estimate treatment effects in the same way we do with real outcomes $\{Y_{i}^{\text{real}}\}_{i=1}^{n}$. This ensures that all ATEs are defined at the individual level, not at the simulation-draw level.
We report sample sizes $n$, number of simulations $K$, and uncertainty measures obtained via a paired bootstrap at the individual level. Specifically, we resample individuals (with replacement) from the fixed analysis sample, e.g., January or December retirees for pensions, or individuals within the 40-year bandwidth for unemployment. For every bootstrap sample, we recompute both the empirical and model ATEs,
$\widehat{\text{ATE}}_{\text{emp}}=\mathbb{E}[Y_{i}^{\text{real}}\mid T_{i}{=}1]\;-\;\mathbb{E}[Y_{i}^{\text{real}}\mid T_{i}{=}0],\qquad\widehat{\text{ATE}}_{\text{model}}=\mathbb{E}[Y_{i}^{\text{sim}}\mid T_{i}{=}1]\;-\;\mathbb{E}[Y_{i}^{\text{sim}}\mid T_{i}{=}0],$
and their difference
$\Delta=\widehat{\text{ATE}}_{\text{model}}-\widehat{\text{ATE}}_{\text{emp}}.$
We use $B=1000$ resamples and form percentile intervals from the empirical distributions. In the tables we therefore report (i) point estimates of $\widehat{\text{ATE}}_{\text{emp}}$ and $\widehat{\text{ATE}}_{\text{model}}$, (ii) their respective 95% confidence intervals, and (iii) the signed discrepancy $\Delta$ with its own confidence intervals.
D.1 Mobility allowance duration